OpenStudy (anonymous):

Suppose f '(2) = 3 and g '(2) = 8 Find h'(2) where h(x) = 5f(x) + 3g(x) +1 h'(2) =

OpenStudy (anonymous):

h'(x) = 5f'(x) +3g'(x) +1 => h'(2) = 5(3) +3(8) +1 =15+24+1 = 40

OpenStudy (anonymous):

is this ur ans!!!

OpenStudy (anonymous):

the answer to the problem said that it was 39:PP its like an online hw assignment

OpenStudy (anonymous):

i dont know how its 39 though

OpenStudy (anonymous):

ohhhhhhhhhhh i'am sorry i made a mistake!!!

OpenStudy (anonymous):

h'(x) = 5f'(x) +3g'(x) +1 => h'(2) = 5(3) +3(8) =15+24 = 39 (if we differentiate 1 we get 0)

OpenStudy (anonymous):

ohhh i see haha thank you:)

OpenStudy (anonymous):

h[x] = 5 f[x] + 3 g[x] + 1 h'[2]=5 f ' [2]+3 g'[2] h'[2]=5*3 + 3*8 h'[2]=15 + 24 h'[2]=39

OpenStudy (anonymous):

myinmi, hello, i think i see why the discrepancy in my problem

OpenStudy (anonymous):

in the problem you have to take the square root of tan^2 theta, but we know that the square root of x^2 is |x|, and for the theta in the problem we have a negative answer